# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ II.4 Other Methods
These schemes are popular in the context of many-body systems and otherwise known under the name of “mean-field approximation” and the “Hartree-approximation” including various generalizations of the latter, (for example, the “Hartree-Fock” method,“ Hrtree-Fock-Bogoliubov” method etc). The basic concept of the scheme consists in approximating the given potential by an exactly solvable one-body-potential. If the original Hamiltonian is H and the approximating Hamiltonian is $H_{0}$, then perhaps the simplest way to ensure self-consistency in the approximation could be imposition of the following constraint :
$<n|H|n>~{}=~{}<n|H_{0}|n>~{},$ (13)
where, the states $|n>$ are the eigenstates of $H_{0}$. In practice, the approximating Hamiltonian may be chosen with a-priori unknown, adjustable parameters, which then get determined through the constraint, eqn.(13) and other requirements such as the variational principle and / or further simplifying conditions. The non-linear feedback that usually characterizes the self-consistency condition can then be easily ensured. Thus, the simple-looking equation, eqn.(13) possesses the potential to include interaction effects non-perturbatively and the simultaneous preservation of the non-linearity of the original Hamiltonian. The above method of implementation often yields results which are reasonably accurate even in the leading order of approximation, which can be further improved through standard recursive procedure as in the case of the Hartree-Scheme.
In the context of ${}\lambda\phi^{4}~{}$ quantum field theory and the QAHO problem, such a scheme was considered in ref.[79] where, however, the expectation value (see,eqn.(13)) was restricted to the ground state only, in implementing the self-consistency requirement. This limitation has been overcome in proposing a “generalised Hartree-method” in ref.[52]. The ‘mean-field’ approach has also been tried in ref.[80].
The main limitations of the method appear to be the lack of uniqueness of the initial choice of the approximating Hamiltonian and the rate of convergence of the recursive procedure for subsequent improvement.
This completes our survey of the various approximation methods in QT that are relevant in the context of the present report. It may be noted that we have described only those schemes which have general applicability. There are however several other methods which are system-specific. These latter often lead to greater accuracy but only at the cost of losing universality of application. Some such schemes will be described in later Sections when we consider specific applications.
@@ A List of Desirable Criteria for an Ideal Scheme of Approximation :
(i)General Applicability and the Universality of the Scheme:
This requirement means that the scheme must be capable of application to general quantum-systems and for arbitrary strength of interactions. Thus, e.g. it should be applicable to quantum mechanics, many-body systems and quantum-field theory both in the non-relativistic and relativistic domains (This criterion is in contrast to “system-specific methods” which are applicable only to the specific system under study and often primarily for achieving high level of accuracy). Additionally it is also required that the scheme should be applicable for arbitrary strength of interaction, thus encompassing both the perturbative- and non-perturbative domains as per standard nomenclature.
@@
The simplicity of formalism and ease of implementation/ computation are obvious desirable features of an ideal AS.
@@
Consistent with the criterion (i) above , reasonable accuracy, i.e. at least to within a few percent of the exact/ numerical results, may be desired for general applications. However, it would be an added virtue if better accuracy can be achieved still retaining the general nature of the scheme.
@@
The scheme may be capable of achieving further improvement in a systematic manner order-by-order over and above a leading order (LO) result and finally
@@
The sequence of higher order corrections in any perturbative framework should either rapidly converge for a convergent sequence or decrease with sufficient rapidity in magnitude, for an asymptotic sequence, in order to be useful for practical implementation.
As stated earlier, it is believed that the above set of criteria may at least form a part of more comprehensive/ exhaustive requirements, if any, on judging a given approximation scheme as an ’ideal’ one. In the next Section we confront the above set of criteria to the extant schemes of approximation in QT.
@@ IV A critique of the various extant schemes of approximation
It is revealed from the above brief, yet hopefully representative , survey (Section-II) that in spite of considerable progress achieved in tackling various problems in interacting quantum systems, an ‘ideal ’ method of approximation based upon the above set of requirements is perhaps yet to be achieved as apparently none of the extant schemes simultaneously satisfy the desired criteria.
In view of the above scenario, there remains a lot of scope to achieve the goal of formulating an ideal scheme based upon the listed criteria. This is then the main motivation for the investigations presented in this report. We elaborate upon this aspect in the following Section.
V Motivation, objectives and the scope of the present report
The present report is devoted to the investigations towards achieving the above ambitious objective, namely, to develop an “ideal” approximation scheme on the stated criteria of general applicability to quantum mechanics and quantum field theory (QFT) and satisfying the criteria of simplicity of formulation, self-consistency; being non-perturbative in nature, yet amenable to systematic improvement in an improved perturbation theory resulting in finally achieving reasonable accuracy.
For simplicity, the scope of the investigations reported here is limited to the quantum physics of the self-interacting Bosonic- systems. In this class, we include the one dimensional quartic anharmonic oscillator (QAHO), quartic double-well oscillator (QDWO), the sextic- anharmonic oscillator (SAHO) and the sextic double-well oscillators (SDWO), the octic-anharmonic oscillator(OAHO) and the $\lambda\phi^{4}$ quantum field theory in physical dimensions and in the massive symmetric phase. These investigations are contained in subsequent Sections and organized as described below:
@@ V.1 Organization of the subsequent sections
We outline the general formulation of the proposed approximation scheme in Section-VI. Hereinafter we refer to this new formulation as “Non-perturbative General Approximation Scheme”(NGAS). In this Section , the non-perturbative nature and the self-consistency of the method, as well as, its general applicability are highlighted. Application to a class of an-harmonic interactions (AHI) in the leading order (LO) is contained in the next Section-VII. In this class, we consider sequentially the cases of the QAHO, the QDWO, the SAHO , the SDWO and the OAHO. The general aspect of the formulation is demonstrated by identical treatment of all these cases within the harmonic approximation. We compute the energy-spectrum and compare the leading order (LO)-results with those obtained from earlier approaches and also with exact numerical results wherever available. In this comparison, it is demonstrated how this simple approach reproduces, in the leading order (LO), the results to within a few percent, of some of the earlier methods employing rather different assumptions and often with sophisticated numerical analysis. In Section-VIII , we demonstrate the flexibility of the present scheme by carrying out the analysis of the AHI by choosing the infinite square-well potential (ISWP) as the input approximation. The stability and flexibility of the scheme is established by reproducing results of comparable accuracy and preservation of the general features of the scheme with this drastically different input.
Next, we extend the formalism to $\lambda\phi^{4}$-quantum field theory in physical dimensions and in the massive symmetric-phase in Section-IX . The renormalization of the bare-parameters of the scheme is carried out and demonstrated to be equivalent to the “Gaussian-effective potential” approach. In particular, the emergence of the non-triviality of the renormalized theory is demonstrated. The study of the properties of the vacuum state of the effective theory in the LO establishes the non-trivial structure of the former. In particular, the equivalence of the present approach to that employing the Bogoliubov-transformations is demonstrated. Further, it is shown that the vacuum of the effective theory corresponds to a condensate of particle-pairs of the free-theory. Moreover,the instability of the perturbative vacuum is established by showing that the vacuum-energy of the effective theory lies much lower than that of the former.
In Section-X we describe the analytic computation of corrections to the LO results in NGAS to arbitrary order in a new formulation of perturbation theory(PT). We refer to this new formulation as the “mean-field perturbation theory (MFPT)” in what follows. In order to compare and contrast the relative merits of the SFPT and MFPT it is deemed necessary to first describe the general features of the two formulations. These are therefore, discussed in Section-XI and Section-XII , respectively for SFPT and MFPT. A common feature in either formulation is the divergence of the perturbation series (PS). Hence it is necessary to elaborate how to make sense of the divergent PS by extracting a finite ”sum” for the former. Hence, we discuss this aspect in some detail also in Section-XI.
The analytic computation of perturbative corrections is achieved by recursive evaluation of the same using as ”tools” , the Feynman-Hellman theorem (FHT) and the hyper-virial theorem (HVT) . In the context of anharmonic interactions and using the standard formulation of perturbation theory(SFPT), these tools were first used by Swenson and Danforth (SD) [81] and it is described succinctly here in Section-XI. The generalization of the same formalism to MFPT is contained in Section- XIII.
In Section-XIV, we describe the application to anharmonic-interactions by computing the perturbation corrections, by extending the application of the SD-method to the case of MFPT via the Feynman-Hellman Theorem (FHT) and the Virial Theorems (VT). The large-order behavior of perturbation-corrections in MFPT establishes the asymptotic nature of the series as expected but demonstrated to be Borel-summable in all cases of the AHI investigated in this report . The Borel-summability of the divergent PS for the case of the QDWO in MFPT is highlighted in Section-XIV as this result is in sharp contrast to the situation in SFPT where Borel summation is frustrated by the “renormalon” -singularity in the Borel-plane. Also contained in this Section-XIV, is the characterization of the leading singularity in the Borel-plane, inferred from the large-order behavior of the perturbation series . In Section-XV the computation of the total perturbative corrections (TPC) to the LO-results are presented and tabulated for the different cases of the AHI by employing two independent methods: (i) by the method of optimal-truncation (MOT) of the asymptotic-series at the term of least magnitude and (ii) by Borel-summation. In the latter case, we evaluate the Borel-Laplace integral by using the technique of conformal mapping for the analytic continuation of the Borel-series beyond its circle of convergence and obtain highly accurate results for the energy of the ground state over a wide range of the coupling strength in each case of the AHI considered herein. In Section-XVI, we conclude with a summary of results presented here and compare and contrast these with those obtained in other investigations in the literature. Possible further applications of NGAS to several other areas are pointed out in Section-XVII.
@@ VI.1 General Formulation for Hamiltonian systems
For simplicity and concreteness, we confine the formulation to one dimensional self -interacting quantum systems with stationary , non-degenerate energy eigenvalues. Consider a generic Hamiltonian $H(g)$ describing such a system where $g$ is the coupling strength of the interaction ( hereinafter we suppress the dependence of $H$ on other variables, for notational simplicity). The first step in the implementation of the scheme consists of choosing an “Approximating Hamiltonian” (AH) denoted by $H_{0}$, which is suitable for the system under investigation and consistent with the basic system-properties such as the symmetries and the given initial- and boundary- conditions for it. Other requirements on the AH are detailed below.
@@ VI.1.1 Prescription of conditions on the approximating Hamiltonian (AH)
The eigenvalue equation to be solved for the system-Hamiltonian is given by:
@@
where we choose the stationary-energy, $E_{n}(g)$ of the quantum-system as the observable. The Hamiltonian, $H(g)$ can be split in to an exactly solvable dominant part $H_{0}$ and a sub-dominant perturbation,
@@
By choice, the AH is defined to be $H_{0}$. Then the implementation of NGAS proceeds by imposing the following conditions on the AH:
(i)The AH, $H_{0}$ is exactly solvable:
This condition requires that eigenvalue equation for $H_{0}$ given by
@@
is exactly solvable, i.e. the eigenstates: $|{\phi}_{n}\rangle$ and the eigenvalues: ${E}_{0}^{n}$ are exactly known. Moreover, the eigenstates are normalized and given by
$\langle\phi_{n}|\phi_{n}\rangle~{}=~{}1$ (15)
We designate this requirement, eqn.(14), as the “condition of exact solvability” (CES).
It is further required that
(ii) $H_{0}$ is chosen to depend upon a set of the free (adjustable) parameters,$\{\alpha_{i}\}$ i.e.
@@
In the next step, we impose the ’condition of equal quantum average’, as described below.
@@ VI.1.2 Condition of equal quantum-average (CEQA)
This condition is defined by the requirement that the AH is constrained to yield the same quantum average (QA) as the original system-Hamiltonian, i.e.
$\langle\phi_{n}|H(g)|\phi_{n}\rangle~{}=~{}\langle\phi_{n}|H_{0}(\{\alpha_{i}\})|\phi_{n}\rangle,$ (16)
for all physical ``$n"$ and ``$g"$ where, $|\phi_{n}\rangle$ is defined by the eigenvalue equation for $H_{0}$, eqn.(14) with ``$n"$ denoting the spectral label. In what follows,eqn.(16) will be referred as the “Constraint of Equal Quantum Average (CEQA)”. The next step is to optimize the approximation as described below:
@@ VI.1.3 Determination of the free parameters in AH by variational-optimization
To determine the parameters { ${}\alpha_{i}~{}$}, we impose CEQA as well as variational optimization of the AH in respect of the still undetermined parameters $(\alpha_{i})$:
$\frac{\partial}{\partial{\alpha_{i}}}\langle H_{0}\rangle=0$ (17)
where, the notation is
$\langle{A}\rangle~{}\equiv~{}\langle\phi_{n}|{A}|\phi_{n}\rangle$ (18)
We refer to this condition, eqn.(17) as the “condition of optimality (CO)”.
This simple procedure, not only completely fixes the free parameters ${\alpha_{i}}$ but also generates the feedback of nonlinearity and $g$-dependence in to the AH. This is elaborated in the next subsection.

### v3

@@ II.4 Other Methods
The self-consistent schemes of approximation are prevalent in the context of many-body systems and are commonly referred to as the ”mean-field approximation” and the ”Hartree-approximation,” including various generalizations such as the ”Hartree-Fock” method and the ”Hartree-Fock-Bogoliubov” method. The fundamental concept of this scheme involves approximating the given potential with an exactly solvable one-body potential. If the original Hamiltonian is denoted as H and the approximating Hamiltonian as $H_{0}$, then a straightforward method to ensure self-consistency in the approximation is to impose the following constraint:
$n|H|n>~=~<n|H_{0}|n>~,$ (13)
where, the states $|n>$ are the eigenstates of $H_{0}$. Practically, the approximating Hamiltonian may be chosen with initially unknown, adjustable parameters, which are subsequently determined through the constraint, eqn.(13), and other criteria such as the variational principle and/or additional simplifying conditions. The non-linear feedback characteristic of the self-consistency condition can thus be readily ensured. Consequently, the ostensibly simple equation, eqn.(13), has the potential to incorporate interaction effects non-perturbatively while preserving the non-linearity of the original Hamiltonian. This method of implementation often yields results that are reasonably accurate even at the leading order of approximation, which can be further refined through a standard recursive procedure, as exemplified by the Hartree-Scheme.
In the context of $\lambda\phi^{4}~$ quantum field theory and the QAHO problem, such a scheme was considered in ref.[79], where, however, the expectation value (see, eqn.(13)) was restricted to the ground state only in implementing the self-consistency requirement. This limitation has been addressed by proposing a ”generalized Hartree-method” in ref.[52]. The ’mean-field’ approach has also been explored in ref.[80].
The primary limitations of the method appear to be the lack of uniqueness in the initial choice of the approximating Hamiltonian and the rate of convergence of the recursive procedure for subsequent improvement.
This concludes our survey of the various approximation methods in QT that are relevant in the context of the present report. It should be noted that we have described only those schemes with general applicability. However, there are several other methods that are system-specific. These latter often lead to greater accuracy but only at the cost of losing universality of application. Some such schemes will be described in later Sections when we consider specific applications.
@@ A List of Desirable Criteria for an Ideal Scheme of Approximation :
(i)General Applicability and the Universality of the Scheme: This criterion necessitates that the scheme be applicable to a wide range of quantum systems and accommodate interactions of arbitrary strength. For instance, it should be relevant to quantum mechanics, many-body systems, and quantum field theory in both non-relativistic and relativistic contexts. This stands in contrast to ”system-specific methods,” which are tailored to particular systems and often prioritize high accuracy. Furthermore, the scheme should be applicable for arbitrary strength of interaction, thereby covering both perturbative and non-perturbative domains as per standard nomenclature.
@@
The formalism’s simplicity and ease of implementation or computation are evidently desirable attributes of an ideal approximation scheme.
@@
In alignment with criterion (i), it is desirable for the scheme to achieve reasonable accuracy, ideally within a few percent of exact or numerical results, for general applications. However, it would be advantageous if greater accuracy could be attained while maintaining the scheme’s general applicability.
@@
The scheme should allow for systematic improvement in an ordered manner, progressing beyond a leading order (LO) result.
@@
In any perturbative framework, the sequence of higher-order corrections should either converge rapidly for a convergent sequence or decrease sufficiently in magnitude as an asymptotic sequence, to ensure practical utility.
As previously mentioned, this set of criteria may constitute part of a more comprehensive set of requirements for evaluating an approximation scheme as ’ideal.’ In the subsequent Section, we will assess these criteria against existing approximation schemes in quantum theory.
@@ IV A critique of the various extant schemes of approximation
The preceding survey (Section-II), though concise, aims to be representative and reveals that despite significant advancements in addressing various challenges within interacting quantum systems, an ’ideal’ approximation method that fulfills the outlined requirements has yet to be realized. It appears that none of the current schemes simultaneously meet the desired criteria. Consequently, there remains substantial potential to develop an ideal scheme based on the specified criteria. This potential constitutes the primary impetus for the investigations detailed in this report. We will further elaborate on this aspect in the following Section.
V Motivation, Objectives, and Scope of the Present Report
This report is dedicated to the pursuit of the aforementioned ambitious objective: to develop an ”ideal” approximation scheme that adheres to the stated objective of broad applicability to quantum mechanics and quantum field theory (QFT). The proposed scheme should meet the criteria of simplicity in its formulation, maintain self-consistency, and exhibit a non-perturbative nature. Additionally, it should allow for systematic enhancement through an improved perturbation theory, ultimately achieving a reasonable degree of accuracy.
To maintain focus on implementation, this study confines its scope to the quantum physics of self-interacting Bosonic systems. The systems under consideration include the one-dimensional quartic anharmonic oscillator (QAHO), quartic double-well oscillator (QDWO), sextic anharmonic oscillator (SAHO), sextic double-well oscillators (SDWO), octic anharmonic oscillator (OAHO), and the $\lambda\phi^{4}$ quantum field theory. These are all analyzed within the context of physical dimensions and specifically within the massive symmetric phase.   These investigations are contained in  subsequent Sections  and organized as described below:
@@ V.1 Organization of the subsequent sections
In Section-VI, we present the general formulation of the proposed approximation scheme, hereafter referred to as the “Non-perturbative General Approximation Scheme” (NGAS). This section emphasizes the non-perturbative nature, self-consistency, and broad applicability of the method. Section-VII addresses the application of this scheme to a class of an-harmonic interactions (AHI) at the leading order (LO). Within this class, we sequentially examine the cases of the QAHO, QDWO, SAHO, SDWO, and OAHO. The formulation’s generality is demonstrated through the uniform treatment of all these cases within the harmonic approximation. We calculate the energy spectrum and compare the leading order (LO) results with those derived from previous approaches and exact numerical results, where available. This comparison illustrates how this straightforward approach reproduces, at the leading order (LO), results that are within a few percent of those obtained by earlier methods, which often involve different assumptions and sophisticated numerical analyses. In Section-VIII, we demonstrate the flexibility of the current scheme by analyzing the AHI using the infinite square-well potential (ISWP) as the input approximation. The stability and flexibility of the scheme are confirmed by reproducing results of comparable accuracy and maintaining the general features of the scheme with this significantly different input. In Section-IX, we advance the formalism to encompass the $\lambda\phi^{4}$-quantum field theory in physical dimensions, concentrating on the massive symmetric phase. The process of renormalization of the bare parameters is executed and demonstrated to align with the ”Gaussian-effective potential” methodology. Importantly, the non-triviality of the renormalized theory is highlighted. An analysis of the vacuum state’s properties within the effective theory at leading order (LO) reveals its non-trivial structure. This approach is shown to be equivalent to the one employing Bogoliubov transformations. Moreover, it is demonstrated that the vacuum of the effective theory is a condensate of particle pairs from the free theory. Additionally, the instability of the perturbative vacuum is confirmed by illustrating that the vacuum energy of the effective theory is considerably lower than that of the perturbative vacuum. In Section-X, we detail the analytic computation of corrections to the leading-order (LO) results in NGAS, extending to arbitrary order through a novel perturbation theory (PT) framework. This innovative framework is termed ”mean-field perturbation theory (MFPT)” in the subsequent discussion. To thoroughly evaluate the relative strengths of the standard formulation of perturbation theory (SFPT) and MFPT, it is crucial to first outline the fundamental characteristics of both approaches. These characteristics are explored in Section-XI and Section-XII, respectively, for SFPT and MFPT. A shared characteristic of both approaches is the divergence of the perturbation series (PS). Therefore, it is essential to clarify the process of interpreting the divergent PS by obtaining a finite ”sum” for the former. This topic is examined in detail in Section-XI. The analytic computation of perturbative corrections is achieved by recursive evaluation of the same using as ”tools” , the Feynman-Hellman theorem (FHT) and the hyper-virial theorem (HVT) . In the context of anharmonic interactions  and using the standard formulation of perturbation theory(SFPT),  these tools were first used by Swenson and Danforth (SD) [81] and  it is described succinctly here in Section-XI. The generalization of the same formalism to MFPT is contained in Section- XIII. In Section-XIV, we elucidate the application to anharmonic interactions by calculating the perturbation corrections, extending the application of the SD-method to the case of MFPT via the Feynman-Hellman Theorem (FHT) and the Virial Theorems (VT). The large-order behavior of perturbation corrections in MFPT confirms the asymptotic nature of the series, as anticipated, yet it is demonstrated to be Borel-summable in all cases of the AHI examined in this report. The Borel summability of the divergent PS for the case of the QDWO in MFPT is emphasized in Section-XIV, as this result starkly contrasts with the situation in SFPT, where Borel summation is impeded by the ”renormalon” singularity in the Borel plane. Additionally, this Section-XIV includes the characterization of the leading singularity in the Borel plane, inferred from the arge-order behavior of the perturbation series. In Section-XV, the computation of the total perturbative corrections (TPC) to the leading-order (LO) results is presented and tabulated for various cases of the AHI, utilizing two independent methodologies: (i) the method of optimal truncation (MOT) of the asymptotic series at the term of least magnitude, and (ii) Borel summation. In the latter approach, the Borel-Laplace integral is evaluated using the technique of conformal mapping for the analytic continuation of the Borel series beyond its circle of convergence, yielding highly accurate results for the ground state energy across a wide range of coupling strengths in each case of the AHI considered herein. In Section-XVI, a summary of the results is provided, with a comparison and contrast to findings from other studies in the literature. Potential further applications of NGAS to several other areas are discussed in Section-XVII.
@@ VI.1 General Formulation for Hamiltonian systems
To ensure clarity and focus, this formulation is limited to one-dimensional self-interacting quantum systems with stationary, non-degenerate energy eigenvalues. We consider a generic Hamiltonian $H(g)$, which characterizes such a system, where $g$ denotes the coupling strength of the interaction. For simplicity in notation, the dependence of $H$ on other variables is omitted. The initial phase of the scheme’s implementation requires the selection of an ”Approximating Hamiltonian” (AH), represented by $H_{0}$. This AH must be well-suited to the system under study and align with essential system properties, including symmetries and the specified initial and boundary conditions. Further criteria for the AH are outlined below.
@@ VI.1.1 Prescription of Conditions on the Approximating Hamiltonian (AH)
The eigenvalue equation to be addressed for the system Hamiltonian is expressed as follows:
@@
where the stationary energy, $E_{n}(g)$, of the quantum system is selected as the observable. The Hamiltonian, $H(g)$, can be decomposed into an exactly solvable dominant component $H_{0}$ and a sub-dominant perturbation,
@@
The term ’approximating Hamiltonian’ is represented as $H_{0}$. The ’New General Approximation Scheme (NGAS)’ is implemented by imposing specific criteria on the AH: (i) The AH, $H_{0}$, must be exactly and analytically solvable. This requirement implies that the eigenvalue equation for $H_{0}$, expressed as
@@
is exactly solvable. In other words, the eigenstates $|{\phi}_{n}\rangle$ and the eigenvalues ${E}_{0}^{n}$ are precisely known. Additionally, the eigenstates are normalized, as indicated by
[…]
This stipulation, encapsulated in equation (14), is referred to as the ”condition of exact solvability” (CES). Furthermore, it is required that (ii) $H_{0}$ is constructed to depend on a set of free (adjustable) parameters, $\{\alpha_{i}\}$, such that
@@
In the subsequent step, the ’condition of equal quantum average’ is applied, as detailed below.
@@ VI.1.2 Condition of equal quantum-average (CEQA)
This condition is defined by the requirement that the AH must yield the same quantum average (QA) as the original system Hamiltonian, as expressed by:
[…]
for all physical n``n" and g``g", where $|\phi_{n}\rangle$ is defined by the eigenvalue equation for $H_{0}$, eqn.(14), with n``n" representing the spectral label. Hereafter, eqn.(16) will be referred to as the “Constraint of Equal Quantum Average (CEQA)”. The subsequent step involves the optimization of the approximation as delineated below:
@@ VI.1.3 Determination of the free parameters in AH by variational-optimization
To determine the parameters { $\alpha_{i}~$}, we implement the CEQA alongwith the variational optimization of the AH with respect to the undetermined parameters $(\alpha_{i})$:
[…]
where the notation is
[…]
We designate this condition, eqn.(17), as the “condition of optimality (CO)”. This methodical approach not only conclusively determines the free parameters $\alpha_{i}$} but also implements the feedback of nonlinearity and $g$-dependence into the AH. This aspect is further discussed in the following subsection.